Honda:

Just check my work for the problem below. my work is in the comments. (refrain from answering until it's up) problem: Using the following equation, find the center and radius of the circle. You must show all work and calculations to receive credit. x2 + 4x + y2 − 6y = −4

Honda:

x^2+4x+y^2-6y=-4 (x^2+4x)+(y^2-6y)=-4 (4/2)^2=4 (6/2)^2=9 (x^2+4x+4)+(y^2-6y+9)=-4+(4+9) (x^2+4x+4)+(y^2-6y+9)=9 (x+2)^2+(y-3)^2=3^2 center=(-2.3) Radius=3

Katrin369:

@honda wrote:
x^2+4x+y^2-6y=-4 (x^2+4x)+(y^2-6y)=-4 (4/2)^2=4 (6/2)^2=9 (x^2+4x+4)+(y^2-6y+9)=-4+(4+9) (x^2+4x+4)+(y^2-6y+9)=9 (x+2)^2+(y-3)^2=3^2 center=(-2.3) Radius=3
You should be good :)

200082741:

@katrin369 You sure it's not 6..?

Katrin369:

Yes I'm sure.

200082741:

@honda put in 3 but if it's wrong try 6

Honda:

@200082741 wrote:
@honda put in 3 but if it's wrong try 6
I only have one chance to get this right.

Katrin369:

Try 3.

200082741:

Oh shoot nvm- yeah it's 3

200082741:

I was solving for x xD mb

Honda:

alright. thanks guys